# A New Era in Computational Pathology: A Survey on Foundation and Vision-Language Models

### v2

@@ Abstract
Recent advances in deep learning have completely transformed the domain of computational pathology (CPath), which in turn altered the diagnostic workflow of pathologists by integrating foundation models (FMs) and vision-language models (VLMs) in their assessment and decision-making process. FMs overcome the limitations of existing deep learning approaches in CPath by learning a representation space that can be adapted to a wide variety of downstream tasks without explicit supervision. VLMs allow pathology reports written in natural language to be used as a rich semantic information source to improve existing models as well as generate predictions in natural language form. In this survey, a holistic and systematic overview of recent innovations in FMs and VLMs in CPath is presented. Furthermore, the tools, datasets and training schemes for these models are summarized in addition to categorizing them into distinct groups. This extensive survey highlights the current trends in CPath and the way it is going to be transformed through FMs and VLMs in the future.
@@
Computational Pathology, Foundation Models, Multi-Modal, Vision-Language Models
@@ I Introduction
In recent years there has been a surge of artificial intelligence (AI) based approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] in CPath owing to wide adoption of digital slide scanners. As a result, large-scale curation and annotation [20, 21] of whole slide images (WSIs) has been made possible which ensured adequate data to train these AI-based models. The goal of these AI-based models is to automate and expedite the diagnosis and prognosis process of CPath. The traditional diagnosis process in CPath is time-consuming and requires experts with extensive domain knowledge. In addition, the wide variation of pathology and heterogeneity between tasks makes it difficult to come up with a unified general approach.
Several research studies have addressed this lack of a unified approach and among the proposed methods, the FMs have gained a lot of attention in recent years[22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. FMs leverage self-supervised learning (SSL) [43] schemes to learn a rich representation in a task-agnostic manner. Owing to self-supervised pre-training (SSPT), FMs do not require large-scale annotated data which is hard to come by in CPath. Furthermore, these models can be trained with a diverse selection of datasets containing tissue samples from different organs and associated with different cancer types, scanner types, etc. Certain research works put a great emphasis on making sure this diversity is present in the pre-training dataset [26, 29]. As a result, the resultant pre-trained model can easily be utilized in a wide range of downstream tasks while maintaining robustness to extreme variation in tissue samples.
Figure 1: Number of publications for FMs and VLMs in pathology (from Google Scholar). The search was done with the keywords “vision-language” + “pathology” for VLMs statistics and “foundation models”+ “pathology” for FMs statistics.
Figure 2: Outline of the major challenges in CPath. Several causes and consequences for each challenge are outlined in addition to how FMs and VLMs address these challenges.
The impact of FMs in CPath can be amplified by integrating the power of VLMs [44] which surged in popularity after the introduction of contrastive language-image pre-training (CLIP) [45] model by OpenAI. Pathology reports, books, educational videos, etc are a rich source of semantic information that can be utilized by VLMs to significantly boost performance which is not possible with vision-only models. When used in conjunction with FMs they can perform like AI pathologists capable of performing a vast array of tasks as evidenced by recent research works [37]. Even though CPath is a specialized field, there is a large increase in the number of publications focusing on FMs and VLMs as shown in Fig. 1 which signals the future direction of this field.
@@ I-A Scope of the Review
    First, articles that focus solely on pathology are included and articles that focus on other areas of the biomedical domain are excluded. As an example, articles like BiomedCLIP [46] with pathology as a subsection of the research are not included in the review.
@@
    Secondly, only vision-language models are included and articles that use other modalities along with vision are excluded. As an example, along with vision, transcriptomics[47] can be used to solve pathological tasks. However, such papers are excluded to maintain the scope.
@@
Figure 3: Visualization of the timeline of recently published work in CPath utilizing FMs and VLMs as well as multi-modal datasets. To maintain transparency we clearly annotate research articles that have been peer-reviewed and articles that are available as pre-print. Furthermore, high-impact pioneering research works published in prominent journals are highlighted. For pre-prints if there are multiple versions, the latest version and the corresponding date are used.
@@ I-B Contribution and Organization
In the past few years, quite a few review articles [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] have been published focusing on CPath. Most of these articles provide a review of the application of deep learning as a whole rather than focusing on a specific subtopic. There are a handful of papers that focus on specific architecture like MIL [18], graph-based models [48, 17], transformer-based models [13, 10], LLMs [19], etc. The contribution of this review article is mentioned below:
@@
    An exhaustive list of multi-modal datasets (Section II) in computational pathology that are being used or can be used in vision-language research is outlined.
@@
    A annotated timeline of the surveyed articles (Fig. 3) which provides a clear idea of the evolution of the FMs and VLMs in CPath. In addition, a project page is available (https://cpath-fms-vlms.github.io/survey/) with additional information about the surveyed articles.
The rest of the article is organized as follows: in section II existing multi-modal datasets are listed along with details about their source, pre-processing techniques, etc. In section III existing FMs in CPath are outlined and descriptions for the vision and vision-language pre-training schemes for these FMs are provided. In section IV an extensive list of VLMs in CPath is provided along with details about their architecture, utilized datasets and contribution. Finally, in section V the paper is concluded.
@@ II Multi-Modal Datasets in Pathology
Figure 4: Different components of multi-modal datasets in computational pathology: Type of datasets, source of data, annotation and pre-processing
@@ II-A Type of Datasets:
The first component is the type of data which can be broadly classified into $55$ categories. Prov-Path [38] fall into the last category which is dissimilar to the rest of the datasets. The dataset for Prov-Path utilizes WSI and the corresponding reports along with histopathology findings, cancer staging, genomic mutation profiles, etc collected by Providence Health System (PHS).
@@
This is in contrast to other datasets in this category like OpenPath [42] which is constructed with Twitter posts and replies and Quilt-1M [52] which is constructed with frames extracted from educational pathology videos. However, the curation and annotation process of this category is much easier as it can automated with hand-crafted algorithms and heuristics. But it comes with the trade-off of noisy data as due to the automated process a lot of artifacts can be present in the data.
The WSI VQA/text category (WSI-VQA [49], PathQABench [37], PathText [57]) contains question and answer pairs or texts associated with WSIs. The most common data source is the cancer genome atlas (TCGA) [20] which contains a large repository of WSI and patient report pairs. The VQA part can be of two types, close-ended question-answer pair and open-ended question-answer pair. Close-ended question-answer pairs are of a multiple-choice type or short-answer type which have pre-defined answers. On the other hand, open-ended question-answer pairs contain answers that are in natural language form. Among the datasets in this category, PathQABench is unique as it contains ROI annotation of WSIs performed by expert pathologists. It has two subsets PathQABench-Public and PathQABench-Private. The former is publicly available as it was constructed with TCGA WSI and reports, and the latter was constructed with in-house data.
@@
The comparison between dataset size is shown in Fig. 5.
@@ II-B Source of Data:
Figure 8: A high-level visualization of different factors of variability in terms of organs, stain types, scanner types, magnification levels, preservation methods, downstream tasks, task levels, pre-training tissue sample sources, etc in foundation models. The research that puts the most emphasis on ensuring variability are RudolfV [26] and PLUTO [29]. Among the tissue sample sources for pre-training TCGA and PAIP are publicly accessible and the rest are proprietary.
@@ III-B Pre-training Workflow and Strategies
In this single diagram, both pre-training strategy for VFM and VLFM is shown. For VFMs, a vision module with some initial weight is leveraged in pre-training. The term “vision module” is used to generally represent a wide range of vision architectures and also modified versions of these architectures with additional layers. The most common architecture is variants of vision image transformers (ViTs) [88] which includes ViT-S, ViT-B, ViT-L, ViT-G and ViT-H. Though there are specialized architectures proposed in BEPH [28] and Prov-GigaPath [38] which uses BEiTv2 [89] and GigaPath architecture respectively. Note that, some models initialize the architectures with ImageNet [90] weights to get an initial vision representation space which can be transformed through pre-training. In Table III, a summary of pre-training strategies in different phases is outlined.
@@ III-B1 Unimodal Vision Pre-training
Among the surveyed articles 3B-CPath [32] and Prov-GigaPath [38] use MAE but in conjunction with DINO and DINOv2 respectively. Research works utilizing iBOT approach include Phikon [33] and CONCH [41].
@@ III-B3 Vision-Language Pre-training
The second widely used approach is the CoCa [97] framework shown in Fig. 14.
@@ III-D Downstream Tasks and Datasets
Cancer Detection [42],[39],[23],[32],[28],[24]
@@
Disease/Cancer/Tissue/Tumor/Molecular SubTyping [42], [41],[39],[38],[23],[31],[28],[26]
@@
Image/Tissue/Tumor/Nuclear Segmentation [41],[31],[26]
@@
Text-to-Image Retrieval [42],[41]
Image-to-Text Retrieval [41]
@@
Mutation Detection/Prediction [38],[31],[32]
@@ III-E Framework, Benchmarking and Adaptation of FMs
PanNuke [42]
@@
WSSS4LUAD [42],[40],[41]
PatchCamelyon (PCam) [27],[26],[24]
MHIST [27],[26],[24]
BACH [28],[31]
SICAP, SICAPv2 [41]
@@
MIDOG [24]
CAMELYON16 [28],[23],[31]
CAMELYON17-WILDS [31],[24]
@@
TCGA CRC-MSI [31],[24]
TCGA-TILs [31],[24]
@@ IV-A Categorization of VLMs
Owing to this, some VLMs follow the FM approach of pre-training the vision or text module. However, as pre-training from scratch requires a large amount of data and computing some VLMs use domain-specific vision modules from HIPT, CTransPath, IBMIL, etc or language modules from PubMedBERT, BioClinicalBERT, etc. This allows the vision or language module to learn an initial vision or language representation space which might not be as rich as pre-training scratch but boosts the performance.

### v3

@@ Abstract
Recent advances in deep learning have completely transformed the domain of computational pathology (CPath). More specifically, it has altered the diagnostic workflow of pathologists by integrating foundation models (FMs) and vision-language models (VLMs) in their assessment and decision-making process. The limitations of existing deep learning approaches in CPath can be overcome by FMs through learning a representation space that can be adapted to a wide variety of downstream tasks without explicit supervision. Deploying VLMs allow pathology reports written in natural language be used as rich semantic information sources to improve existing models as well as generate predictions in natural language form. In this survey, a holistic and systematic overview of recent innovations in FMs and VLMs in CPath is presented. Furthermore, the tools, datasets and training schemes for these models are summarized in addition to categorizing them into distinct groups. This extensive survey highlights the current trends in CPath and its possible revolution through the use of FMs and VLMs in the future.
@@
Computational pathology, foundation models, multi-modal, vision-language models.
@@ I Introduction
In recent years there has been a surge of artificial intelligence (AI)-based approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] in computational pathology (CPath) owing to wide adoption of digital slide scanners. In fact, large-scale curation and annotation [20, 21] of whole slide images (WSIs) has ensured adequate data for training of these AI-based models. The goal of generating such AI-based models is to automate and expedite the diagnosis and prognosis process of CPath. Traditional diagnosis process in CPath is time-consuming and requires experts with extensive domain knowledge. In addition, the wide range of cancer types and grades as well as the heterogeneity among diagnosis/prognosis tasks, makes it challenging to come up with a unified general approach.
Several research studies have addressed the issue with a unified approach and among the proposed methods, foundation models (FMs) have gained a lot of attention in recent years[22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. More specifically, FMs leverage self-supervised learning (SSL) [43] schemes to learn a rich representation in a task-agnostic manner. Owing to self-supervised pre-training (SSPT), FMs do not require large-scale annotated data which is hard to come by in CPath. Furthermore, these models can be trained with a diverse selection of datasets containing tissue samples from different organs and associated with different cancer types, as well as scanner types. Certain research works focus on making sure this diversity is present in the pre-training dataset [26, 29]. As a result, the resultant pre-trained model can easily be utilized in a wide range of downstream tasks while maintaining robustness to extreme variation in tissue samples.
Figure 1: Number of publications for FMs and VLMs in pathology (from Google Scholar). The search keywords include “vision-language” + “pathology” for VLMs statistics and “foundation models”+ “pathology” for FMs statistics.
Figure 2: Outline of major challenges in CPath. Several causes and consequences for each challenge are outlined in addition to how FMs and VLMs address these challenges.
The impact of FMs in CPath can be amplified by integrating the power of vision-language models (VLMs)  [44]. More specifically, VLMs have surged in popularity after the introduction of contrastive language-image pre-training (CLIP) model by OpenAI [45]. Pathology reports, books, educational videos, etc. are rich sources of semantic information that can be utilized by VLMs to significantly boost the performance. This is not typically possible with vision-only models. When deployed in conjunction with FMs, they can perform as AI-based pathologists capable of performing a vast array of tasks as evidenced by recent literature [37]. Even though CPath is a specialized area of study in the medical field, there is a large increase in the number of publications focusing on FMs and VLMs in pathology as shown in Fig. 1 demonstrating a dominant future direction of this field.
@@ I-A Scope of the Review
    First, articles that focus solely on pathology are included. So, and articles with a focus on other areas of the biomedical domain are excluded. As an example, articles like BiomedCLIP [46] with pathology as a subsection of the research are not included in this review paper.
@@
    Secondly, only vision-language models are included and articles that use other modalities along with vision are excluded. As an example, along with vision, transcriptomics[47] can be used to address pathology tasks. However, such papers are excluded to maintain the scope.
@@
Figure 3: Visualization of the timeline of recently published work in CPath utilizing FMs and VLMs as well as multi-modal datasets. To maintain transparency, we clearly annotate research articles that have been peer-reviewed and articles that are available as pre-prints. Furthermore, high-impact pioneering research works published in prominent journals are highlighted. For pre-prints if there are multiple versions, the latest version and the corresponding date are used.
@@ I-B Contribution and Organization
In the past few years, quite a few review articles [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] have been published focusing on CPath. Most of these articles provide a review of the application of deep learning as a whole rather than focusing on a specific subtopic. There are handful of papers that focus on specific architecture like MIL [18], graph-based models [48, 17], transformer-based models [13, 10], LLMs [19], etc. The contributions of this review article are summarized below:
@@
    An exhaustive list of multi-modal datasets (Section II) in CPath that are being used or can be used in vision-language research are outlined.
@@
    Annotated timeline of the surveyed articles (Fig. 3) which provides a clear idea of the evolution of the FMs and VLMs in CPath. In addition, a project page is available (https://cpath-fms-vlms.github.io/survey/) with additional information about the surveyed articles.
The rest of the article is organized as follows: in section II existing multi-modal datasets are listed along with details about their source, pre-processing techniques, etc. Section III outlines existing FMs in CPath and descriptions for the vision and vision-language pre-training schemes for those FMs are provided. In section IV an extensive list of VLMs in CPath is presented along with details of their architectures, utilized datasets and contributions. Finally, section V concludes the paper.
@@ II Multi-Modal Datasets in Pathology
Figure 4: Different components of multi-modal datasets in computational pathology: Type of datasets, sources of data, annotation and pre-processing
@@ II-A Type of Datasets:
The first component is the type of data which can be broadly classified into $55$ categories. Prov-Path [38] fall into the last category which is dissimilar to the rest of the datasets. The dataset for Prov-Path utilizes WSI and the corresponding reports along with histopathology findings, cancer staging, genomics mutation profiles, etc. collected by Providence Health System (PHS).
@@
This is in contrast to other datasets in this category like OpenPath [42] which is constructed with Twitter posts and replies and Quilt-1M [52] which is constructed with frames extracted from educational pathology videos. However, the curation and annotation process of this category is much easier as it can be automated with hand-crafted algorithms and heuristics. However, it comes with an inevitable noisy data as due to the automated process a lot of artifacts can be present in the data.
The WSI VQA/text category (WSI-VQA [49], PathQABench [37], PathText [57]) contains question and answer pairs or texts associated with WSIs. The most common data source is the cancer genome atlas (TCGA) [20] which contains a large repository of WSI and patient report pairs. The VQA part can be of two types, close-ended question-answer pair and open-ended question-answer pair. Close-ended question-answer pairs are of a multiple-choice type or short-answer type which have pre-defined answers. On the other hand, open-ended question-answer pairs contain answers that are in natural language form. Among the datasets in this category, PathQABench is unique as it contains region of interest (ROI) annotation of WSIs performed by expert pathologists. It has two subsets as PathQABench-Public and PathQABench-Private. The former is publicly available as it was constructed with TCGA WSIs and reports, and the latter was constructed with in-house data.
@@
A comparison of dataset sizes is shown in Fig. 5.
@@ II-B Source of Data:
Figure 8: A high-level visualization of different factors of variability in terms of organs, stain types, scanner types, magnification levels, preservation methods, downstream tasks, task levels, pre-training tissue sample sources, etc in foundation models. The research that puts the most emphasis on ensuring variability are RudolfV [26] and PLUTO [29]. Among the tissue sample sources for pre-training, TCGA and PAIP are publicly accessible and the rest are proprietary.
@@ III-B Pre-training Workflow and Strategies
In this single diagram, both pre-training strategy for VFM and VLFM is shown. For VFMs, a vision module with some initial weight is leveraged in pre-training. The term “vision module” is used to generally represent a wide range of vision architectures and also modified versions of these architectures with additional layers. The most common architecture is variants of vision image transformers (ViTs) [88] which includes ViT-S, ViT-B, ViT-L, ViT-G and ViT-H. Though there are specialized architectures proposed in BEPH [28] and Prov-GigaPath [38] which uses BEiTv2 [89] and GigaPath architecture, respectively. Note that, some models initialize the architectures with ImageNet [90] weights to get an initial vision representation space which can be transformed through pre-training. In Table III, a summary of pre-training strategies in different phases is outlined.
@@ III-B1 Unimodal Vision Pre-training
Among the surveyed articles 3B-CPath [32] and Prov-GigaPath [38] use MAE but in conjunction with DINO and DINOv2, respectively. Research works utilizing iBOT approach include Phikon [33] and CONCH [41].
@@ III-B3 Vision-Language Pre-training
The second widely used approach is the contrastive captioner (CoCa) [97] framework shown in Fig. 14.
@@ III-D Downstream Tasks and Datasets
Disease/Cancer/Tissue/Tumor/Molecular SubTyping [42], [41],[39],[38],[23],[31], [28],[26],[30],[35],[29]
Cancer Detection [22],[42],[39],[23],[32],[28],[24],[35]
@@
Image/Tissue/Tumor/Gland/Nuclei Segmentation [41],[31],[26], [35],[29]
@@
Text-to-Image Retrieval [42],[41],[30]
Image-to-Text Retrieval [41],[30]
@@
Mutation Detection/Prediction [38],[31],[32],[33]
@@ III-E Framework, Benchmarking and Adaptation of FMs
PanNuke [42],[29]
HEST-1K [22]
DLBCL-Morph [22]
TissueNet [35]
GlaS [29]
CRAG [35]
RenalCell [30]
SkinCancer [30]
UBC-OCEAN [36]
Chaoyang [36]
@@
WSSS4LUAD [42],[40],[41],[30],[36]
PatchCamelyon (PCam) [22],[27],[26],[24],[36]
MHIST [22],[27],[26],[24]
BACH [28],[31],[30],[36]
SICAP, SICAPv2 [41], [30]
@@
MIDOG [22], [24],[35]
CAMELYON16 [28],[23],[31],[36],[33],[35]
CAMELYON17-WILDS [22],[31],[24],[29],[36],[33]
@@
TCGA CRC-MSI [31],[36],[24]
TCGA-TILs [22], [24],[31]
@@ IV-A Categorization of VLMs
Owing to this, some VLMs follow the FM approach of pre-training the vision or text module. However, as pre-training from scratch requires a large amount of data and computing some VLMs use domain-specific vision modules from HIPT, CTransPath, IBMIL, etc or language modules from PubMedBERT, BioClinicalBERT, etc. This allows the vision or language module to learn an initial vision or language representation space which might not be as rich as pre-training scratch but boosts the performance. A summary of the choice of encoder type and pre-training strategy is provided in Table IX.